# SkyScenes: A Synthetic Dataset for Aerial Scene Understanding

### v1

@@ 3.2 SkyScenes Image Generation
We generate SkyScenes images from Carla [7] by taking the previously mentioned considerations into account. Curating images from Carla [7] broadly consists of two key steps: (1) positioning the agent camera in an aerial perspective and (2) procedurally guiding the agent within the scene to capture images. We accomplish the first by mimicking a UAV perspective in Carla [7] by positioning the ego vehicle (with RGB, semantic and depth sensors) based on specified (high) altitude () and pitch () values to generate aerial views (see Fig. 3).444This also requires setting other scene – weather, daytime, etc. – and camera – notably the (field of view) and (image resolution) – parameters. Once positioned, the agent is translated by fixed amounts to traverse the scene and capture images from various viewpoints (detailed in Algo. 1 in the appendix). Initially, we generate datapoints for each of the town variations under ClearNoon conditions using the baseline setting. Subsequently, following
the traversal algorithm (see Algo. 1 in the appendix),
@@
Avoiding Overly Correlated Frames for Viewpoints.
Carla [7] uses a traffic manager with a PID controller to control the egocentric vehicle based on current pose, speed, and a list of waypoints at every pre-defined time step. Curating images at every time step (or tick) results in highly correlated frames with little change in object position. Since overly correlated frames are not very useful when training models for static scene understanding, we move the camera by a fixed distance multiple times before saving a frame. This also helps with moving dynamic actors by a considerable amount in the scene. Additionally, pedestrian objects are regenerated before saving an image, which adds randomness to the spawning and placement of pedestrians and further reduces correlation between frames.
| Layout | RS (%) | HS (%) |
| 1 Town01 | ||
| 2 Town02 | ||
| 3 Town03 | ||
| 4 Town04 | ||
| 5 Town05 | ||
| 6 Town06 | ||
| 7 Town07 | ||
| 8 Town10HD | ||
| 9 All | ||
| 10 UAVid [19] |
| mIoU | |||
| Eval Data | HS | human | All |
| 1 S | |||
| 2 S | ✓ | 61.79 | 84.07 |
| 3 SU | |||
| 4 SU | ✓ | 10.21 | 47.09 |
Adequate Representation of humans. Real-world scenes often exhibit a long-tailed distribution in pixel proportions, particularly in aerial images where variations in object sizes and camera positions contribute to significant under-representation of the tail classes (in Fig. 2, for the shared set of classes across UAVid [19] (aerial) and Cityscapes [5] (ground), we can see that the class distributions are different and aerial images are significantly more heavy tailed). As a result, naively spawning humans (rarest class) in Carla [7] is detrimental for eventual task performance – for the human class, a SkyScenes trained DAFormer [11](with HRDA [12] source training; MiT-B5 [36] backbone) model leads to an in-distribution performance of mIoU and out-of-distribution (SkyScenes UAVid [19]) performance of mIoU. To counter this under-representation issue, we design an algorithm, HumanSpawn()555More details in Sec. B.1 of the appendix., to explicitly spawn more human instances while curating SkyScenes images. HumanSpawn() increases human instances by per snapshot, improving the proportion of densely annotated humans in SkyScenes by approximately times (see Table. 3 & Fig. 4). This improvement in human representation is also evident in eventual task performance, with in-distribution and out-of-distribution mIoUs for humans increasing from to () and to () respectively (see Table. 3.)
@@ 3.3 SkyScenes: Dataset Details
Annotations. We provide semantic, instance and depth annotations for every image in SkyScenes. Semantic annotations in SkyScenes by default are across classes. These are building, fence, pedestrian, pole, roadline (markings on road), road, sidewalk, vegetation, cars, wall, traffic sign, sky, bridge, railtrack, guardrail, traffic light, water, terrain, rider, bicycle, motorcycle, bus, truck and others (see Fig. 1 for an example).666We provide detailed definitions in Sec. B.2 of the appendix.
Training, Validation and Test Splits. SkyScenes has images per town (across towns) for each of the weather and daytime conditions, and height & pitch combinations, resulting in a total of images. We use % ( images) of the dataset for training models, with % ( images) each for validation and testing. While creating train, val and test splits, we collect equal number of samples from each town by dividing each town-specific traversal sequence into segments: the initial % for training, the next % for testing, and the final % of the segment for validation. Moreover, within each split, we ensure that every viewpoint is accompanied by its different variations across weather, daytime, height, and pitch settings. This safeguards against any potential cross-contamination across different splits while ensuring fair representation and equal distributions of all variations.
Class Distribution(s). In Fig. 6, we compare the number of densely annotated pixels in SkyScenes with those in Syndrone [25] (another synthetic dataset) and UAVid [19] (a real aerial dataset). Compared to both the real and prior synthetic counterparts, we show that SkyScenes is specifically curated to ensure adequate tail class representation (more human and vehicle pixels) to facilitate model learning. Additionally, in Fig. 5, we highlight how the distribution of classes changes across variations within SkyScenes– rural and urban map layouts and height and pitch specifications. SkyScenes exhibits substantial diversity in class-distributions across such conditions, allowing these individual conditions to serve as diagnostic splits to assess model sensitivity (see Sec. 4.2)
@@ 4 Experiments
We conduct semantic segmentation experiments with SkyScenes to assess a few different factors. First, we check if training on SkyScenes is beneficial for real-world transfer. Second, we check if SkyScenes can augment real-world training data in low and full shot regimes. Third, we check if variations in SkyScenes can be used to assess sensitivity of trained models to changing conditions. Finally, we check if using additional modality information (depth) can help improve aerial scene understanding.
Synthetic and Real Datasets. We compare real-world generalization performance of training on SkyScenes with SynDrone [25], a recently proposed synthetic aerial dataset also curated from Carla [7]. We assess performance on real-world aerial datasets – UAVid [19], AeroScapes [22], ICG Drone [14]. Since different datasets have different class vocabularies and definitions, for our experiments, we adapt the class vocabulary of the synthetic source dataset to that of the target real-world datasets.777We detail class merging and assignment schemes used for these experiments in Table. 8, Table. 9 and Table. 10 in the appendix. Additionally, since different real aerial datasets have been captured from different heights and pitch angles, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions. We provide additional details for the real aligned synthetic data selection and model evaluation in Sec. D.1 of the appendix.
Models. We use both CNN – DeepLabv2 [3] (ResNet-101 [10]) – and transformer – DAFormer [11] (with HRDA [12] source training; MiT-B5 [36] backbone) – based semantic segmentation architectures for our experiments. We provide implementation details surrounding our experiments in Sec. C of the appendix.
@@
| 4 SynDrone | |||
| 5 SkyScenes | 47.09 | 40.72 | 25.91 |
| 6 Train-on-Target | |||
| Source | (Target) Real-World mIoU () | |||||
| UAVid | AeroScapes | ICG Drone | ||||
| vehicle | human | vehicle | person | vehicle | person | |
| 1 SynDrone | ||||||
| 2 SkyScenes | 63.64 | 10.21 | 80.99 | 3.09 | 39.71 | 45.89 |
| 3 Train-on-Target | ||||||
@@ 4.1 SkyScenes Real Transfer
SkyScenes trained models generalize well to real-settings. In Table. 4, we show how models trained on SkyScenes exhibit strong out-of-the box generalization performance on multiple real world datasets. We find that SkyScenes pretraining exhibits stronger generalization compared to SynDrone [25] across both CNN and transformer segmentation backbones. In Table. 5, we show how generalization improvements are more pronounced for under-represented tail classes (vehicles and humans)888Model performance comparisons for all the classes are provided in Table 11, Table. 12 and Table. 13 in the appendix.
SkyScenes can augment real training data. In addition to zero-shot real-world generalization, akin to other synthetic aerial datasets, we also show how SkyScenes is useful as additional training data when labeled real-world data is available. In Fig. 7, for SkyScenes UAVid [19], we compare models trained only using of the UAVid [19] training images with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time. We find that in low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance. We find this to be especially beneficial for under-represented classes in aerial imagery (such as humans and vehicles). We discuss more such results across other real-world datasets in Sec. D.2 of the appendix.
| Train | Test mIoU () | ||
| Clear | Cloudy | Rainy | |
| 1 Clear | |||
| 2 Cloudy | |||
| 3 Rainy | |||
| Train | Test mIoU () | ||
| Noon | Sunset | Night | |
| 1 Noon | |||
| 2 Sunset | |||
| 3 Night | |||
| Train | Test mIoU () | |
| Rural | Urban | |
| 1 Rural | ||
| 2 Urban | ||
| Test mIoU () | ||||
| Height | Pitch | |||
| 1 m | 48.50 | |||
| 2 m | ||||
| 3 m | ||||
@@ 4.2 SkyScenes as a Diagnostic Framework
As noted earlier, the images we curate in SkyScenes contain several variations – ranging from different weather and daytime conditions, rural and urban map layouts, and different height and pitch combinations (see Fig. 5 for variations in class distributions). We curate images under such diverse conditions in a controlled manner – ensuring the same spatial coordinates for variations, same spatial coordinates and settings across different weather and daytime conditions, same number of images across layouts. This allows us to assess the sensitivity of trained models to one factor of variation (, , daytime, weather, map layout) by changing that specific aspect. We summarize some takeaways from such experiments in Table. 6.
In Table. 6 (a), we show how models trained in a certain weather condition are best at generalizing to the same condition at test-time. We make similar observations for daytime variations in Table. 6 (b). In Table. 6 (c), we show how models trained in rural conditions fail to perform well in urban test-time conditions and vice-versa. In Table. 6 (d), we evaluate a model trained under moderate conditions under different variations. We find that as altitudes increase, trained models are better at recognizing objects from oblique () viewpoints. We provide exhaustive quantitative comparisons in Sec. D.3 of the appendix.
| Sensors | SkyScenes Test mIoU () | |||||||
| clutter | building | road | tree | low-veg. | vehicle | human | Avg | |
| 1 RGB | ||||||||
| 2 RGB+D | 90.64 | 95.97 | 94.87 | 89.41 | 74.36 | 86.87 | 50.47 | 83.22 |
4.3 SkyScenes Enables Multi-modal Recognition
Sensors on UAVs in deployable settings are not limited to RGB cameras. It is common to have UAVs deployed in the real-world with additional modality sensors, such as depth. Additional sensor modalities can also potentially help improve aerial scene understanding. In Table. 7, we check if assisting RGB with Depth observations for SkyScenes viewpoints can help improve aerial semantic segmentation using M3L [20], a model capable of multimodal segmentation. Similar to our DAFormer [11] experiments, we consider a SegFormer equivalent version of M3L [20] (with an MiT-B5 [36] backbone). We test RGB and RGB+D models trained under (moderate viewpoint) conditions on SkyScenes and find that incorporating additional Depth observations can substantially improve recognition performance. This demonstrates that annotated images in SkyScenes can be used to train multimodal scene-recognition models.
@@ 5 Conclusion
We introduce SkyScenes, a large-scale densely-annotated dataset of synthetic aerial scenes curated from unmanned aerial vehicle (UAV) perspectives. We collect SkyScenes images from Carla by first aerially situating an agent (to get an aerial perspective) and then procedurally tele-operating the agent through the scene to capture aerial frames with corresponding semantic, instance and depth annotations. Our careful curation process ensures that SkyScenes images are carefully curated across diverse weather, daytime, map, height and pitch conditions, with accompanying metadata that enables reproducing the same viewpoint (spatial coordinates and perspective) under differing conditions.
Through our experiments, we demonstrate how (1) SkyScenes trained models can generalize to real-world settings, (2) SkyScenes images can augment labeled real-world data in low-shot regimes, (3) SkyScenes can serve as a diagnostic framework to assess model sensitivity to changing conditions and (4) additional sensors, such as Depth, in SkyScenes can facilitate development of multi-modal aerial scene understanding models.
Lastly, we plan on updating SkyScenes with evolving considerations for real-world aerial scene-understanding – improved realism, additional anticipated edge cases – as more and more features are supported in the underlying simulator. We intend to release both the dataset and associated generation code for SkyScenes publicly, and hope that our experimental findings encourage further research using SkyScenes for aerial scenes.
@@
Appendix A Appendix
This appendix is organized as follows. In Sec. B, we provide more details on different aspects of dataset – including the procedural image curation algorithm, algorithm to ensure appropriate representation of human instances and brief descriptions of all the classes in SkyScenes. Then, we describe experimental details in Sec. C – class-merging schemes used for our SkyScenes Real transfer experiments, train / val / test splits, experimental details for SkyScenes diagnostic setup to probe model vulnerabilities. Sec. D provides more quantitative and qualitative experimental results.
Appendix B SkyScenes Details
B.1 Image Generation Algorithms
We curate SkyScenes using Carla [7] simulator. The generation process broadly consists of two key steps: (1) positioning the agent camera in an aerial perspective and (2) procedurally guiding the agent within the scene to capture images. We accomplish the first by mimicking a UAV perspective in Carla [7] by positioning the ego vehicle (with RGB, semantic, depth and instance segmentation sensors) based on specified altitude () and pitch () values to generate aerial views. This also requires setting other scene information like – town, weather, and daytime. The camera (field of view) and (image resolution) are also set.
To maintain the adequate representation of tail classes, especially for humans we incorporate structured spawning of humans using Algo. 2. Carla [7] has a limit on the number of actors that can be spawned in a scene, which depends on factors such as the type and size of the town, number of lanes to spawn vehicles, and amount of sidewalk area. To overcome this limitation, we decided to bring the actors into the field of view of the camera instead of having them spread out in the scene. We developed an algorithm to find all the points to spawn pedestrians in the field of view using the camera location and spawn them like vehicles on roads. After taking a snapshot, we destroy the spawned pedestrians and repeat the process. Manual spawning not only increases the number of human instances and their proportion but also aligns their placement with real-world settings. The steps involved in manual spawning instances of humans are summarized below:

### v2

@@ 3.2 SkyScenes Image Generation
We generate SkyScenes images from Carla [10] by taking the previously mentioned considerations into account. Curating images from Carla [10] broadly consists of two key steps: (1) positioning the agent camera in an aerial perspective and (2) procedurally guiding the agent within the scene to capture images. We accomplish the first by mimicking a UAV perspective in Carla [10] by positioning the ego vehicle (with RGB, semantic and depth sensors) based on specified (high) altitude () and pitch () values to generate aerial views (see Fig. 2(a)).555This also requires setting other scenes – weather, daytime, etc. – and camera (notably the (field of view) and image resolution x) parameters. Once positioned, the agent is translated by fixed amounts to traverse the scene and capture images from various viewpoints (detailed in Sec. 0.B.1 in the appendix). Initially, we generate data points for each of the town variations under ClearNoon conditions using the baseline m setting.
Subsequently, following
the traversal algorithm (see Sec. 0.B.1 in the appendix),
@@
| mIoU | ||
| Eval Data | human | All |
| SkyScenes | ||
| 1 Before HS | ||
| 2 After HS | 61.79 | 84.07 |
| SkyScenes UAVid | ||
| 3 Before HS | ||
| 4 After HS | 10.21 | 47.09 |
@@
Avoiding Overly Correlated Frames for Viewpoints. Carla [10] uses a traffic manager with a PID controller to control the egocentric vehicle based on current pose, speed, and a list of waypoints at every pre-defined time step. Curating images at every time step (or tick) results in highly correlated frames with little change in object position. Since overly correlated frames are not very useful when training models for static scene understanding, we move the camera by a fixed distance multiple times before saving a frame. This also helps with moving dynamic actors by a considerable amount in the scene. Additionally, pedestrian objects are regenerated before saving an image, which adds randomness to the spawning and placement of pedestrians, reducing the correlation between frames.
Adequate Representation of humans. Real-world scenes often exhibit a long-tailed distribution in pixel proportions, particularly in aerial images where variations in object sizes and camera positions contribute to significant under-representation of the tail classes (in Fig. 2, for the shared set of classes across UAVid [27] (aerial) and Cityscapes [8] (ground), we can see that the class distributions are different and aerial images are significantly more heavy-tailed). As a result, naively spawning humans (rarest class) in Carla [10] is detrimental for eventual task performance – for the human class, a SkyScenes trained DAFormer [16] (with HRDA [17] source training; MiT-B5 [55] backbone) model leads to an in-distribution performance of mIoU and out-of-distribution (SkyScenes UAVid [27]) performance of mIoU. To counter this under-representation issue, we design an algorithm, HumanSpawn() (see Sec. 0.B.1 in appendix), to explicitly spawn more human instances while curating SkyScenes images. HumanSpawn() increases human instances by per snapshot, improving the proportion of densely annotated humans in SkyScenes by approximately times (see Fig. 3 (a) & Fig. 3 (b)). This improvement in human representation is also evident in eventual task performance, with in-distribution and out-of-distribution mIoUs for humans increasing from to () and to () respectively (see Table. 3 (c)).
@@ 3.3 SkyScenes: Dataset Details
Annotations. We provide semantic, instance and depth annotations for every image in SkyScenes. Semantic annotations in SkyScenes by default are across classes. These are building, fence, pedestrian, pole, roadline (markings on road), road, sidewalk, vegetation, cars, wall, traffic sign, sky, bridge, railtrack, guardrail, traffic light, water, terrain, rider, bicycle, motorcycle, bus, truck and others (see Fig. 1 for an example and Sec. 0.B.2 in appendix for definitions)
Training, Validation and Test Splits. SkyScenes has images per town (across towns) for each of the weather and daytime conditions, and height & pitch combinations, resulting in a total of images. We use % ( images) of the dataset for training models, with % ( images) each for validation and testing (see Sec. 0.C.2 in appendix).
Class Distribution(s). In Fig. 4, we highlight how the distribution of classes changes across variations within SkyScenes– rural and urban map layouts and height and pitch specifications. SkyScenes exhibits substantial diversity in class distributions across such conditions, allowing these individual conditions to serve as diagnostic splits to assess model sensitivity (see Sec. 4.2).
@@ 4 Experiments
| 3 SynDrone | |||
| 4 SkyScenes | 47.09 | 40.72 | 25.91 |
| Rein (DINOv2) | |||
| 5 SynDrone | 54.92 | ||
| 6 SkyScenes | 43.96 | 28.10 | |
| Source | (Target) Real-World IoU () | |||||||
| UAVid | AeroScapes | ICG Drone | ||||||
| vehicle | human | vehicle | person | vehicle | person | |||
| DAFormer (MiT-B5) [17] | ||||||||
| 1 SynDrone | ||||||||
| 2 SkyScenes | 63.64 | 10.21 | 80.99 | 3.09 | 39.71 | 45.89 | ||
| Rein (DINOv2) [54] | ||||||||
| 3 SynDrone | ||||||||
| 4 SkyScenes | 75.14 | 25.52 | 87.71 | 21.67 | 50.91 | 77.93 | ||
We conduct semantic segmentation experiments with SkyScenes to assess a few different factors. First, we check if training on SkyScenes is beneficial for real-world transfer. Second, we check if SkyScenes can augment real-world training data in low and full shot regimes. Third, we check if variations in SkyScenes can be used to assess the sensitivity of trained models to changing conditions. Finally, we check if using additional modality information (depth) can help improve aerial scene understanding.
Synthetic and Real Datasets. We compare real-world generalization performance of training on SkyScenes with SynDrone [42], a recently proposed synthetic aerial dataset also curated from Carla [10] featuring different conditions across different map layouts. We assess performance on real-world aerial datasets – UAVid [27], AeroScapes [34], ICG Drone [20]. Since different datasets have different class vocabularies and definitions, for our experiments, we adapt the class vocabulary of the synthetic source dataset to that of the target real-world datasets (see Sec. 0.C.1 in appendix for class merging and assignment schemes). Additionally, since different real aerial datasets have been captured from different heights and pitch angles, we train models on subsets of synthetic datasets that are aligned with corresponding real data conditions. We provide additional details for the real aligned synthetic data selection and model evaluation in Sec. 0.D.3 in the appendix.
Models. We use (1) CNN – DeepLabv2 [4] (ResNet-101 [14]), (2) transformer – DAFormer [16] (with HRDA [17] source training; MiT-B5 [55] backbone) and (3) Vision Foundation Model – Rein [54] (LoRA [19] fine-tuned Dino-V2 [36] backbone) based semantic segmentation architectures for our experiments. We provide implementation details for our experiments in Sec. 0.C in appendix.
Figure 5: SkyScenes RGB Images and corresponding depth images generated using depth sensor for , , and ClearNoon setting across four different town layouts.
Sensors
SkyScenes Test IoU ()
clutter
building
road
tree
low-veg.
vehicle
human
Avg
1 RGB
2 RGB+D
90.64
95.97
94.87
89.41
74.36
86.87
50.47
83.22
Table 5: Multi-modal Segmentation in SkyScenes. We evaluate M3L [28] multimodal segmentation architectures with MiT-B5 [55] backbones using RGB and RGB+D data in SkyScenes. Additional sensors improve aerial scene understanding significantly across various classes in UAVid [27].
@@ 4.1 SkyScenes Real Transfer
SkyScenes trained models generalize well to real-settings. In Table. 3, we show how models trained on SkyScenes exhibit strong out-of-the box generalization performance on multiple real world datasets. We find that SkyScenes pretraining exhibits stronger generalization compared to SynDrone
[42] across both CNN and transformer segmentation backbones. In Table. 3, we show how generalization improvements are more pronounced for under-represented tail classes (vehicles and humans). Comparison across all classes is provided in Tables 11, 12 and 13 in appendix.
SkyScenes can augment real training data. In addition to zero-shot real-world generalization, akin to other synthetic aerial datasets, we also show how SkyScenes is useful as additional training data when labeled real-world data is available. In Table. 4, for SkyScenes UAVid [27], we compare models trained only using , , of the UAVid [27] training images with counterparts that were either pretrained using SkyScenes data or additionally supplemented with SkyScenes data at training time. We find that in low-shot regimes (when little “real” world data is available), SkyScenes data (either explicitly via joint training or implicitly via finetuning) is beneficial in improving recognition performance (see Sec. 0.D.5 of appendix).
@@ 4.2 SkyScenes as a Diagnostic Framework
As noted earlier, the images we curate in SkyScenes contain several variations – ranging from different weather and daytime conditions, rural and urban map layouts, and different height and pitch combinations (see Fig. 4 for variations in class distributions). We curate images under such diverse conditions in a controlled manner – ensuring the same spatial coordinates for variations, same spatial coordinates and settings across different weather and daytime conditions, the same number of images across layouts.
This allows us to assess the sensitivity of trained models to one factor of variation (, , daytime, weather, map layout) by changing that specific aspect. We summarize some takeaways from such experiments in Table. 6.
In Table. 6 (a), we show how models trained in a certain weather condition are best at generalizing to the same condition at test-time. We make similar observations for daytime variations in Table. 6 (b). In Table. 6 (c), we show how models trained in rural conditions fail to perform well in urban test-time conditions and vice-versa. In Table. 6 (d), we evaluate a model trained under moderate m conditions under different variations. We find that as altitudes increase, trained models are better at recognizing objects from oblique () viewpoints. We provide exhaustive quantitative comparisons in Sec. 20 in the appendix.
4.3 SkyScenes Enables Multi-modal Dense Prediction
Sensors on UAVs in deployable settings often include modalities beyond RGB cameras, such as depth sensors. These additional modalities can significantly enhance aerial scene understanding. In Table. 5, we investigate the impact of augmenting RGB data with depth observations from SkyScenes viewpoints on aerial semantic segmentation using M3L [28], a multimodal segmentation model. Similar to our DAFormer [16] experiments, we consider a SegFormer equivalent version of M3L [28] (with an MiT-B5 [55] backbone). We test RGB and RGB+D models trained under (moderate viewpoint) conditions on SkyScenes and find that incorporating additional Depth observations can substantially improve recognition performance. This demonstrates that images in SkyScenes can be used to train multimodal scene-recognition models.
@@ 5 Conclusion
We introduce SkyScenes, a large-scale densely-annotated dataset of synthetic aerial scene images curated from unmanned aerial vehicle (UAV) perspectives. SkyScenes images are generated using Carla by situating an agent aerially and procedurally tele-operating it through the scene to capture frames with semantic, instance, and depth annotations. Our careful curation process ensures that SkyScenes images span across diverse weather, daytime, map, height, and pitch conditions, with accompanying metadata that enables reproducing the same viewpoint (spatial coordinates and perspective) under differing conditions.
Through our experiments, we demonstrate that: (1) SkyScenes-trained models generalize well to real-world settings, (2) SkyScenes augments labeled real-world data in low-shot scenarios, (3) SkyScenes serves as a diagnostic tool for assessing model sensitivity to varied conditions, and (4) incorporating additional sensors like depth enhances multi-modal aerial scene understanding.
We aim to enhance SkyScenes with improved realism, additional anticipated edge cases, and support for 3D perception tasks aligning with advancements in our simulator (additional details in Sec. 0.F of appendix) We have publicly released the dataset and associated generation code and hope that our experimental findings encourage further research using SkyScenes for aerial scenes.
@@
Appendix 0.A Overview
The appendix is organized as follows. In Sec. 0.B.1, we provide more details on different aspects of dataset – including the procedural image curation algorithm, algorithm to ensure appropriate representation of human instances, brief descriptions of all the classes in SkyScenes and comparisons between data distribution of SkyScenes and real datasets. Then, we describe experimental details in Sec. 0.C – class-merging schemes used for our SkyScenes Real transfer experiments, train / val / test splits and experimental details for SkyScenes diagnostic setup to probe model vulnerabilities. Sec. 0.D provides more quantitative and qualitative experimental results.
Appendix 0.B SkyScenes Details
0.B.1 Image Generation Algorithms
We curate SkyScenes using Carla [10] simulator. The generation process broadly consists of two key steps: (1) positioning the agent camera in an aerial perspective and (2) procedurally guiding the agent within the scene to capture images. We accomplish the first by mimicking a UAV perspective in Carla [10] by positioning the ego vehicle (with RGB, semantic, depth and instance segmentation sensors) based on specified altitude () and pitch () values to generate aerial views. This also requires setting other scene information like – town, weather, and daytime. The camera (field of view) and = (image resolution) are also set.
To maintain the adequate representation of tail classes, especially for humans we incorporate structured spawning of humans using Algo. 2. Carla [10] has a limit on the number of actors that can be spawned in a scene, which depends on factors such as the type and size of the town, number of lanes to spawn vehicles, and amount of sidewalk area. To overcome this limitation, we decided to bring the actors into the field of view of the camera instead of having them spread out in the scene. We developed an algorithm to find all the points to spawn pedestrians in the field of view using the camera location and spawn them like vehicles on roads. After taking a snapshot, we destroy the spawned pedestrians and repeat the process. Manual spawning not only increases the number of human instances and their proportion but also aligns their placement with real-world settings.
As SkyScenes images are curated by teleoperating over entire maps (rural or urban) across multiple layouts that differ substantially in class distributions (Fig. 4 left in the paper). Since these frames are stored with corresponding geographical (layout identifier) and positional (spatial location) metadata, filtering data splits to avoid overlap across physical layout regions is always possible.
The steps involved in manual spawning instances of humans are summarized below: